OpenStudy (anonymous):

Please help I have no idea how to do this IN 2005, the total waste generated in a certain country was 6.801 *10^11 pounds. Also, in 2005 the country population was 4.01*10^8 people. Determine the garbage per person in that country in the year 2005.

OpenStudy (anonymous):

garbage per person \[=\frac{ 6.801*10^{11} }{ 4.01*10^{8} }=?\]

OpenStudy (anonymous):

so thats how i find it?

OpenStudy (mathmale):

You're looking for a "rate." If you drive a car, you observe your rate of distance covered over time all the time; it's right there on your dashboard speedometer. Surjithayer has tried to show you the "rate" involved in this problem by typing out the following: 6.801 * 10^11 pounds of garbage per 4.01 * 10^8 people. What are the units of measurement here? pounds per person. That's a rate. Reduce the following: \[\frac{ 6.801*10^{11}pounds.of.garbage }{ 4.01*10^{8^{}}people }\] All you have to do now is to reduce this fraction as much as possible. Divide 4.01 into 6.801 and write the result with 3 decimal place accuracy. Reduce \[\frac{ 10^{11} }{ 10^{8} }\] appropriately. Then you'll be done. Be certain to include units of measurement in your response.

OpenStudy (anonymous):

when i evaluated it it came to 1696.00997506234 ?

OpenStudy (anonymous):

Thanks you! c:

OpenStudy (mathmale):

Again, please write 6.801/4.01 with 2 (not 3) decimal points. You could write the result as 1.70, and then multiply that by 10^3, and then provide the units of measurement.